Find the measure of each angle in triangle
It is an 180 degree
A is x
B is 3x
C is 8x+6

Answer by Alan3354(69443)

It is an 180 degree
A is x
B is 3x
C is 8x+6
-------------
The sum of the angles of all triangles is 180º.
x + 3x + 8x+6 = 180
12x + 6 = 180
12x = 174
x = 14.5º
3x = 43.5º
8x+6 = 122º